OpenStudy (anonymous):

Solve the system by elimination.  -2x+2y+3z=0  -2x-y+z=-3  2x+3y+3z=5 

OpenStudy (anonymous):

@StudyGurl14 by any chance do you know this?

OpenStudy (studygurl14):

Ooh...yuck. I know how, but I really hate systems of equations with three variables...

OpenStudy (studygurl14):

Okay, so do you know what it means by "elimination"?

OpenStudy (anonymous):

Im guessing I have to eliminate variables or numbers or something?

OpenStudy (studygurl14):

sorta.

OpenStudy (studygurl14):

-2x+2y+3z=0 -2x-y+z=-3 2x+3y+3z=5

OpenStudy (anonymous):

Okay, what does it mean?

OpenStudy (studygurl14):

Okay, notice how all the equations start with either 2x or -2x?

OpenStudy (studygurl14):

It is easier to show you, rather than tell you

OpenStudy (anonymous):

yes I see

OpenStudy (studygurl14):

You can eliminate the x variables of the equations by adding/subtracting the equations from each other. take the first equation and the third equation -2x+2y+3z=0 2x+3y+3z=5 ______________ 0 + 5y + 6 z = 5

OpenStudy (anonymous):

okay so do that with them all? like 1,2 then 1,3 then 2,3?

OpenStudy (studygurl14):

And then, you can do the same with the first two equations -2x-y+z=-3 multiply by -1 2x + y - z = 3 Next, -2x+2y+3z=0 2x + y - z = 3 _______________ 0 + 3y + 2z = 3

OpenStudy (anonymous):

what do i do after that?

OpenStudy (studygurl14):

Now you have these two equations: 5y + 6z = 5 3y + 2z = 3

OpenStudy (studygurl14):

You want to eliminate one more variable so you can solve for the other variable.

OpenStudy (anonymous):

okay..

OpenStudy (studygurl14):

3y + 2z = 3 multiply by -3 -9y - 6z = -9 See how it is changed?

OpenStudy (anonymous):

yes but why multiply by -3? and what would multiply the 5y + 6z = 5 by would it be -5?

OpenStudy (studygurl14):

5y + 6z = 5 -9y - 6z = -9 _______________ -4y + 0 = -4 Now what?

OpenStudy (studygurl14):

I multiplied it by -3 so that the 2z would change to -6z, which, when added to the 6z from the other equation, equals zero. Make sense?

OpenStudy (anonymous):

yes! So only one equation needs that not both?

OpenStudy (studygurl14):

right. in some cases, you might need to multiply both equations by something, but not in this case

OpenStudy (studygurl14):

So, you have -4y = -4 What does y = ?

OpenStudy (anonymous):

Oh okay.. oh um.. lol uhhh -4?

OpenStudy (studygurl14):

Divide both sides by -4. What is -4/-4?

OpenStudy (anonymous):

1?

OpenStudy (studygurl14):

Correct! So...y = 1

OpenStudy (anonymous):

thats it? or is there more?

OpenStudy (studygurl14):

There's more...lol. You still have to find the values for x and z

OpenStudy (anonymous):

oh lol darn it.... how do i do that?

OpenStudy (studygurl14):

Okay, so we have this equation from before: 5y + 6z = 5 Now that we know y =1, how do we solve for z?

OpenStudy (anonymous):

Im not sure..

OpenStudy (anonymous):

maybe do the same thing or something similar?

OpenStudy (studygurl14):

Plug in 1 for y, and solve for z 5(1) + 6z = 5 5 + 6z = 5 6z = 0 z = ?

OpenStudy (anonymous):

would it be 6 divided by 6?

OpenStudy (anonymous):

1?

OpenStudy (studygurl14):

Nope. Hint: Any number times 0 = 0

OpenStudy (anonymous):

oh 6 x 0? so 0?

OpenStudy (studygurl14):

Correct. So z = 0

OpenStudy (studygurl14):

So we know z = 0, and y = 1. How do we find the value of x?

OpenStudy (anonymous):

would we use multiplication again?

OpenStudy (studygurl14):

we would plug in again.

OpenStudy (anonymous):

using multiplication?

OpenStudy (studygurl14):

Let's use the very first original equation. -2x+2y+3z=0

OpenStudy (studygurl14):

If z = 0, and y= 1, how would we find x? Using this equation ^

OpenStudy (anonymous):

plug it in: -2(1)+2y+3(0)=0 ?

OpenStudy (studygurl14):

Close. We don't know the value of x yet. We know y = 1, so you would replace y with 1, not x

OpenStudy (anonymous):

Oh i mean -2x+2y(1)+3(0)=0 haha sorry

OpenStudy (anonymous):

Yeah I mistyped lol

OpenStudy (studygurl14):

Right!

OpenStudy (studygurl14):

Okay, so you have: -2x+2(1)+3(0)=0 What now?

OpenStudy (anonymous):

man.. um.. i think like this? -2x+2(1)+3(0)=0 -2x+2+3=0 after that Im a bit stuck but i am starting to understand!

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

lol 5!

OpenStudy (studygurl14):

actually..no

OpenStudy (studygurl14):

3(0) = ?

OpenStudy (anonymous):

3? lol unsure now :P

OpenStudy (studygurl14):

Remember this rule. Whenever a number is being multiplied by 0, the result is 0. Whenever a number (except for 0) is being multiplied by 1, the answer is that number. So, in this situation, is 3 being multiplied by 1 or 0?

OpenStudy (anonymous):

oh sorry the answer is 0 wow im dumb

OpenStudy (studygurl14):

Yep. And no you're not. It is easy to forget the simple stuff once you start dealing with super hard stuff that occupies all the space in your brian..lol

OpenStudy (studygurl14):

Okay, so back to -2x+2(1)+3(0)=0 This would simplify to... -2x + 2 + 0 = 0 What's the next step?

OpenStudy (anonymous):

lol yeah my "brian" hurts right now hahaahah :P

OpenStudy (studygurl14):

lol.

OpenStudy (anonymous):

-2x + 2?

OpenStudy (studygurl14):

Yep...I named my brain Brian. lol

OpenStudy (studygurl14):

Yep -2x + 2 = 0 What now?

OpenStudy (anonymous):

HAHA!! Wait how does 2x + 2 equal zero? :/

OpenStudy (anonymous):

@StudyGurl14 ? :(

OpenStudy (studygurl14):

Because. That was the original equation...go back a few posts and look. You can't eliminate the equals sign...or the zero on the other side. Only the one zero on the left side of the = sign

OpenStudy (anonymous):

oh okay so what would be the next step?

OpenStudy (anonymous):

@StudyGurl14 why do you keep leaving meee we are almost done!!

OpenStudy (studygurl14):

sorry i have to check on my own quesiton

OpenStudy (studygurl14):

ok. so next step...

OpenStudy (anonymous):

oh I thought you were mad at me lol

OpenStudy (studygurl14):

-2x + 2 = 0 - 2 -2 _______________ -2x = -2 What does x equal? And of course I'm not mad at you.

OpenStudy (anonymous):

4? lol thank god i need you!

OpenStudy (studygurl14):

Nope.... Divide both sides by -2 What is -2/-2?

OpenStudy (anonymous):

oh i multiplied! its 1 :)

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

correct

OpenStudy (studygurl14):

So.... x = 1 y = 1 z = 0

OpenStudy (anonymous):

So what do i do with those? am I done or do I need to do something else? @StudyGurl14

OpenStudy (studygurl14):

You're done

OpenStudy (studygurl14):

or should I say "I'm done" lol, jk. xD

OpenStudy (anonymous):

No you should you helped so much thanks again Einstein!! :D

OpenStudy (studygurl14):

lol, you're welcome.

OpenStudy (studygurl14):

Hey, have you written me a fan testimonial yet? *hint hint* lol, xD

OpenStudy (anonymous):

No way I thought I did!

OpenStudy (studygurl14):

sometimes it glitches. maybe try again?

OpenStudy (anonymous):

Yeah I refreshed.. so weird this will be my third time writing it!!

OpenStudy (anonymous):

I wrote one :)

OpenStudy (studygurl14):

Thanks so much!

OpenStudy (studygurl14):

what you wrote was so nice :)

OpenStudy (anonymous):

Of course I really value your help!!! hey for the first and third equation did you divide to get 0 + 3y + 2z = 3 @StudyGurl14

OpenStudy (studygurl14):

no I added

OpenStudy (anonymous):

added them and the other two to get the final two equations?